# Fine-grained Spatial-Temporal Perception for Gas Leak Segmentation

### v1

@@ 1 Introduction
Gas leaks are unintended events that can occur due to damaged or faulty connections, corrosion, or improper installation. Whether it is natural gas, methane, or other hazardous gases, their escape into the atmosphere can lead to explosions, fires, and long-term health complications. Detecting gas leaks is essential for early intervention and mitigation of potential disasters. Traditional leak detection methods involve inspectors manually using instruments near potential leak areas to assess the status of the pipeline. These approaches are inefficient and hazardous to inspectors due to the potentially harmful gas. In contrast, video surveillance systems can provide accurate and real-time monitoring without human labor. Thus, designing a video processing algorithm to recognize leaks by analyzing motion patterns and shape variations is a better option for gas leak segmentation. Such algorithms can process continuous video streams in real-time, providing faster and safer detection compared to manual inspections.
@@
Gas leaks possess characteristics distinct from common objects. They are colorless, odorless, and invisible to human senses. Although infrared cameras can visualize leaks [1], distinguishing them often remains challenging because of their non-rigid shape and blurry texture. In many cases, segmentation is only possible when the leak is in motion. For example, Lu et al. [2] proposes Gaussian-based background modeling for adaptive gas leak segmentation. However, this method faces significant limitations in industrial scenarios due to the irregular morphology of gas and the variability of the background. Common motion extraction method based on background subtraction and optical flow [3] fails due to the variable shape of the object and severe noise disturbances.
@@
Recently, deep-learning techniques have shown robustness in object segmentation [4]. Fan et al. [5] first located coarse concealed objects and then refined them in the decoder. Wang et al. [6], based on YOLOv7 [7], demonstrates the capabilities in detecting visible non-rigid objects like dust. For those objects in RGB images, methods relying solely on spatial information may be effective. However, segmenting blurry objects in infrared cameras becomes challenging if models neglect motion information. Yan et al. [8] incorporates optical flow to learn motion features, and Yang et al. [3] employs unsupervised learning to generate pixel-level motion labels in CNNs. However, both approaches are susceptible to noise interference. Memory-based segmentation methods [9, 10, 11] struggle with the variety of gas leak shapes. Moreover, redundant memory significantly impacts segmentation performance. Cheng et al. [12] integrates boundary optimization and optical flow for motion analysis, showing potential for gas leak segmentation. However, it struggles to distinguish objects which are similar to leaks in texture, like clouds. To the best of our knowledge, few methods can effectively segment blurry and non-rigid objects, like gas leaks, in infrared videos.
@@
Fig. 1: The architecture of FGSTP. The model processes every current frame along with an adjacent frame each time. Each current frame needs to be processed twice with two adjacent frames. The encoder denoises the input frame and extracts the multi-scale features, while the decoder optimizes boundaries through GRA and NCD modules. $f_{i}^{j}$ denotes the feature of the ith frame and jth scale. The Consecutive Temporal Correlation (CTC) block captures motion information. $f^{j}$ denotes the CTC output in the jth scale. The Fine-Grained Spatial Perception (FSP) module refines spatial features.
@@
In this paper, we address these challenges by proposing a fine-grained spatial-temporal network for non-rigid, blurry object segmentation in infrared video. We adopt a multi-scale transformer-based [13] encoder to mitigate video noise. To extract representative motion features of gas leaks, we refine relevant motion clues across multiple scales using a proposed Consecutive Temporal Correlation module (CTC). We also propose the Fine-grained Spatial Perception module (FSP) to capture spatial features by combining the motion features with encoder features and distilling informative semantic features of leaks. Besides, boundary determination is critical yet challenging for polymorphic objects, as their lack of clear boundaries makes them blend with the background. Our decoder can optimize boundaries from coarse prediction to accurate mask. We conduct experiments on the GasVid [1] dataset, for which we manually annotate gas leak regions frame by frame to ensure the accurate representation of leak area and effective model training. In summary, our contributions include: (1) We propose an end-to-end architecture that is specifically to segment non-rigid and blurry objects in infrared videos; (2) We design CTC and FSP modules to obtain key motion clues and perceive shapes or boundary features from encoder output.
@@ 2.1 Framework Overview
Fig.1 shows the architecture of FGSTP. The model consists of four components: (1) Transformer-based encoder denoises and obtains multi-scale features; (2) Consecutive Temporal Correlation (CTC) captures key motion; (3) Fine-grained Spatial Perception (FSP) refines spatial features through the residual connection between encoder features and CTC outputs; (4) The decoder optimizes predicted masks boundaries. FGSTP takes the inference frame with two adjacent frames as input to identify the location and shape of the target object. The model output is a binary mask of the inference frame.
FGSTP is a one-stage model and requires no additional auxiliary inputs. It is designed for segmenting blurry non-rigid objects like gas leaks. The shape of leaks change significantly in the short term. Our research finds that the two adjacent frames contain sufficient information for the inference frame in this kind of video object segmentation. Therefore our model just calculates 3 frames (two adjacent and one inference frames) each time. This operation not only reduces redundant computations but also helps FGSTP capture inter-frame differences more prominently.
@@ 2.2 FGSTP Architecture
(1) Encoder. The Pyramid Vision Transformer (PVT) [13] is the model backbone to extract the global features from three input frames. The encoder has four layers, which generate global features at different-scales. The size of features in each layer is $C\times H/2^{i+1}\times W/2^{i+1}$, where $i\in\{1,2,3,4\}$, $H,W,C$ are height, width and channels. Previous work finds that the last three layers reserve more high-level semantic features with reduced dimensions [12]. Therefore we put the $2nd,3rd,$ and $4th$ layers features into the multi-scale network to extract local features $f_{t}^{i}$ ($i\in\{2,3,4\}$).
@@ 3.2 Metrics and Configurations
Unlike normal object segmentation methods, we don’t use any pre-trained backbone or model due to the specific characteristics of our dataset. Pre-trained models on RGB datasets may negatively impact our experiments. For a fair comparison, we train our model and benchmark models from scratch. The input frames are resized to 352$\times$352. We train for 60 epochs with learning rate $1\times 10^{-4}$. All models are trained on an NVIDIA RTX 4090 GPU.
@@ 3.3 Results
Table 1: Quantitative comparison on various benchmarks in GasVid dataset.
@@
Fig. 4: Visualization results in GasVid Dataset. Our model prediction is the most accurate in different situations. $i.e$ close (camera) distance with complex background (cloud or birds interference, 1467), long distance with complex background (1476), medium distance with clear background (2559), long distance with clear background (2563), close distance with clear background (2566)
@@
We also provide visualized results in Fig.4. Due to MG’s failure and space limitations, we select the results of the best five models in all test videos. Our model consistently produces the most accurate masks across various scenarios. For Video-1467 (first row), the background contains clouds and birds, with the clouds having appearances similar to gas leaks. In these challenging cases, only our model correctly segments the leak. Video-1476 is the most difficult sample because of its long camera distance and cloud-covered background. All models, except ours, generate masks with incorrect shapes and locations. Video-2559 also involves a long camera distance but lacks clouds. SLT-Net correctly localizes the leak in this video, but unlike our model, it fails to accurately match its shape. Videos 2559 and 2566 are relatively easier and all models successfully segment leaks; however, our model’s predictions have more details and ours are the most similar to the ground truth (GT). These results demonstrate that our model not only captures more details in simpler scenarios but also shows more stable and accurate performance in challenging conditions.
@@ 3.4 Ablation Studies
Table 2: Ablation studies results on CTC and FSP modules of FGSTP in GasVid dataset
@@ 4 Conclusion
This paper presents an innovative method for gas leak segmentation that achieves better accuracy and adaptability across diverse scenarios. Our approach incorporates an encoder to reduce noise, a temporal module to refine motion clues, and a spatial perception to effectively distinguish target object. Experimental results demonstrate that our model outperforms SOTA benchmarks. Additionally, ablation studies confirm the critical role of our two core modules for robust performance. Finally, visualized results highlight that our model not only captures finer details in easier cases but also handles various challenging situations effectively.

### v2

@@ 1 Introduction
Gas leaks are unintended events that can occur due to damaged or faulty connections, corrosion, or improper installation. Whether it is natural gas, methane, or other hazardous gases, their escape into the atmosphere can lead to explosions, fires, and long-term health complications. Detecting gas leaks is essential for early intervention and mitigation of potential disasters. Traditional leak detection methods involve inspectors manually using instruments near potential leak areas to assess the status of the pipeline. These approaches are inefficient and hazardous to inspectors due to the potentially harmful gas. In contrast, video surveillance systems can provide accurate and real-time monitoring without human labor. Thus, designing a video processing algorithm to recognize leaks by analyzing motion patterns and shape variations is a better option for gas leak segmentation. Such algorithms can process continuous video streams in real time, providing faster and safer detection compared to manual inspections.
@@
Gas leaks possess characteristics distinct from common objects. They are colorless, odorless, and invisible to human senses. Although infrared cameras can visualize leaks [12], distinguishing them often remains challenging due to their non-rigid shape and blurry texture. In many cases, segmentation is only possible when the leak is in motion. For example, Lu et al. [8] proposed Gaussian-based background modeling for adaptive gas leak segmentation. However, this method faces significant limitations in industrial scenarios due to the irregular morphology of gas and the variability of the background. Common motion extraction method based on background subtraction and optical flow [16] fails due to the variable shape of the object and severe noise disturbances.
@@
Recently, deep learning techniques have shown robustness in object segmentation [18]. Fan et al. [6] first located coarse concealed objects and then refined them in the decoder. Wang et al. [13], based on YOLOv7 [11], demonstrated the capabilities in detecting visible non-rigid objects like dust. For those objects in RGB images, methods relying solely on spatial information may be effective. However, segmenting blurry objects in infrared cameras becomes challenging if models neglect motion information. Yan et al. [15] incorporated optical flow to learn motion features, and Yang et al. [16] employed unsupervised learning to generate pixel-level motion labels in CNNs. However, both approaches are susceptible to noise interference. Memory-based segmentation methods [1, 17, 2] struggle with the variety of gas leak shapes. Moreover, redundant memory significantly impacts segmentation performance. Cheng et al. [3] integrated boundary optimization and optical flow for motion analysis, showing potential for gas leak segmentation. However, it struggles to distinguish objects that are similar to leaks in texture, such as clouds. To the best of our knowledge, few methods can effectively segment blurry and non-rigid objects, like gas leaks, in infrared videos.
@@
Fig. 1: The architecture of FGSTP. The model processes every current frame along with an adjacent frame each time. Each current frame needs to be processed twice with two adjacent frames. The encoder denoises the input frame and extracts the multi-scale features, while the decoder optimizes boundaries through the GRA and NCD modules. $f_{i}^{j}$ denotes the feature of the ith frame and jth scale. The Consecutive Temporal Correlation (CTC) block captures motion information. $f^{j}$ denotes the CTC output in the jth scale. The Fine-grained Spatial Perception (FSP) module refines spatial features.
@@
In this paper, we address these challenges by proposing a fine-grained spatial-temporal network for non-rigid, blurry object segmentation in infrared video. We adopt a multi-scale transformer-based [14] encoder to mitigate video noise. To extract representative motion features of gas leaks, we refine relevant motion clues across multiple scales using a proposed Consecutive Temporal Correlation module (CTC). We also propose the Fine-grained Spatial Perception module (FSP) to capture spatial features by combining the motion features with encoder features and distilling informative semantic features of leaks. Besides, boundary determination is critical yet challenging for polymorphic objects, as their lack of clear boundaries makes them blend with the background. Our decoder can optimize boundaries from coarse prediction to an accurate mask. We conduct experiments on the GasVid [12] dataset, for which we manually annotate gas leak regions frame by frame to ensure the accurate representation of leak area and effective model training. In summary, our contributions include: (1) We propose an end-to-end architecture that is specifically to segment non-rigid and blurry objects in infrared videos; (2) We design CTC and FSP modules to obtain key motion clues and perceive shapes or boundary features from encoder output.
@@ 2.1 Framework Overview
Fig.1 shows the architecture of FGSTP. The model consists of four components: (1) Transformer-based encoder denoises and obtains multi-scale features; (2) Consecutive Temporal Correlation (CTC) captures key motion; (3) Fine-grained Spatial Perception (FSP) refines spatial features through the residual connection between encoder features and CTC outputs; (4) The decoder optimizes predicted mask boundaries. FGSTP takes the inference frame with two adjacent frames as input to identify the location and shape of the target object. The model output is a binary mask of the inference frame.
FGSTP is a one-stage model and requires no additional auxiliary inputs. It is designed for segmenting blurry non-rigid objects like gas leaks. The shape of leaks changes significantly in the short term. Our research finds that the two adjacent frames contain sufficient information for the inference frame in this kind of video object segmentation. Therefore our model calculates 3 frames (two adjacent and one inference frame) each time. This operation not only reduces redundant computations but also helps FGSTP capture inter-frame differences more prominently.
@@ 2.2 FGSTP Architecture
(1) Encoder. The Pyramid Vision Transformer (PVT) [14] is the model backbone to extract the global features from three input frames. The encoder has four layers, which generate global features at different scales. The size of features in each layer is $C\times H/2^{i+1}\times W/2^{i+1}$, where $i\in\{1,2,3,4\}$, $H,W,C$ are height, width and channels. Previous work finds that the last three layers reserve more high-level semantic features with reduced dimensions [3]. Therefore we put the $2nd,3rd,$ and $4th$ layers features into the multi-scale network to extract local features $f_{t}^{i}$ ($i\in\{2,3,4\}$).
@@ 3.2 Metrics and Configurations
Unlike normal object segmentation methods, we don’t use any pre-trained backbone or model due to the specific characteristics of our dataset. Pre-trained models on RGB datasets may negatively impact our experiments. For a fair comparison, we train our model and benchmark models from scratch. The input frames are resized to $\times352$. We train for 60 epochs with learning rate $1\times 10^{-4}$. All models are trained on an NVIDIA RTX 4090 GPU.
@@ 3.3 Results
Table 1: Quantitative comparison of various benchmarks on GasVid dataset.
@@
Fig. 4: Visualization results on GasVid Dataset. Our model prediction is the most accurate in different situations, $i.e.,$ close (camera) distance with complex background (cloud or birds interference, 1467), long distance with complex background (1476), medium distance with clear background (2559), long distance with clear background (2563), close distance with clear background (2566)
@@
We also provide visualized results in Fig.4. Due to MG’s failure and space limitations, we select the results of the best five models on all test videos. Our model consistently produces the most accurate masks across various scenarios. For Video-1467 (first row), the background contains clouds and birds, with the clouds having appearances similar to gas leaks. In these challenging cases, only our model correctly segments the leak. Video-1476 is the most difficult sample because of its long camera distance and cloud-covered background. All models, except ours, generate masks with incorrect shapes and locations. Video-2559 also involves a long camera distance but lacks clouds. SLT-Net correctly localizes the leak in this video, but unlike our model, it fails to accurately match its shape. Videos 2559 and 2566 are relatively easier and all models successfully segment leaks; however, our model’s predictions have more details and ours are the most similar to the ground truth (GT). These results demonstrate that our model not only captures more details in simpler scenarios but also shows more stable and accurate performance in challenging conditions.
@@ 3.4 Ablation Studies
Table 2: Ablation studies results of CTC and FSP modules in FGSTP on GasVid dataset
@@ 4 Conclusion
This paper presents an innovative method for gas leak segmentation that achieves better accuracy and adaptability across diverse scenarios. Our approach incorporates an encoder to reduce noise, a temporal module to refine motion clues, and a spatial perception to effectively distinguish the target object. The experimental results demonstrate that our model outperforms the SOTA benchmarks. Additionally, ablation studies confirm the critical role of our two core modules for robust performance. Finally, visualized results highlight that our model not only captures finer details in easier cases but also handles various challenging situations effectively.